OpenStudy (anonymous):

plz help will give medal and be fan. Evaluate if x = 2.5 and y = –14.6. x – |y|

OpenStudy (mathstudent55):

Substitute the values you were given for x and y in the expression: 2.5 - |-14.6| Do you know how to take the absolute value of -14.6?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

A. –17.1 B. 17.1 C. 12.1 D. –12.1

OpenStudy (anonymous):

can i just get the answer plz

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

Think of absolute value of a number as the distance from that number to zero on the number line. It can never be negative. For example: |0| = 0 |1| = 1 |-1| = 1 |5| = 5 |-7| = 7

OpenStudy (anonymous):

i have more mr selena can you help me with more :)

OpenStudy (mathstudent55):

2.5 - |-14.6| = 2.5 - 14.6 = -12.1

OpenStudy (anonymous):

Provide the reason for Step 4 in the following proof. Prove: –(a +b) + a = –b

OpenStudy (anonymous):

A. Identity Property of Addition B. Commutative Property of Addition C. Definition of Subtraction D. Additive Inverse Property

OpenStudy (mathstudent55):

@Ms.Selena gave you a wrong answer and with no explanation. Keep asking questions and getting no explanations and wrong answers and you'll do really well.

OpenStudy (anonymous):

Dude he just wanted the answer, and btw I got the same answer as you so basically you're saying you got it wrong too.

OpenStudy (anonymous):

@jukeboxhero I'm sorry I don't understand the second!

OpenStudy (anonymous):

o ok umm one sec

OpenStudy (anonymous):

Provide the reason for Step 4 in the following proof. Prove: –(a +b) + a = –b 7bchart.gif A. Identity Property of Addition B. Commutative Property of Addition C. Definition of Subtraction D. Additive Inverse Property

OpenStudy (anonymous):

let me get the chart

OpenStudy (mathstudent55):

@Ms.Selena 1. If for you -12.1 and 12.1 are the same, you have no business "helping" people with math questions. 2. Helping means helping the person understand the probelm, not just throwing an answer.

OpenStudy (anonymous):

D is not C you are mistaking by that

OpenStudy (anonymous):

mistaken

OpenStudy (mathstudent55):

On the last step, a and zero were added together, and the result is a. Adding zero to a number does not change the number. A number after having zero added to it keeps its "identity." Is this enough of a clue?

OpenStudy (mathstudent55):

@Ms.Selena You are correct. I stand corrected. Your answer was correct. D is -12.1 which is correct. I misread it above. Still I don't think that just answering a question with no explanation is helpful to the poster. He didn't learn anything. It's as if he had copied his homework from yours.